# Multi-wavelength Emission from Jets and Magnetically Arrested Disks in Nearby Radio Galaxies: Application to M87

### v1

@@ Appendix A Magnetization parameter for power-law tail distribution
$\dot{N}_{e,\rm inj}=\dot{N}_{0}\begin{cases}\left(\frac{\gamma}{4\sigma}\right)^{-p}\ (\gamma<4\sigma)\\ \left(\frac{\gamma}{4\sigma}\right)^{-p_{\rm tail}}\ (4\sigma<\gamma<\gamma_{\rm rad})\\ \end{cases}.$ (A1)
@@
$L_{\rm rec}$ $\approx 2l_{\rm rec}^{2}\frac{B_{\rm rec}^{2}}{8\pi}\beta_{\rm rec}c$ (A2)
$\simeq 7.9\times 10^{41}M_{9}\dot{m}_{-4}\beta_{\rm rec,-1}\Phi_{\rm rec,1.2}^{2}\ \rm erg\ s^{-1}.$ (A3)
@@
$L_{\rm max}=\frac{L_{\rm rec}}{\frac{1}{2-p}\left[1-\left(\frac{E_{\sigma}}{m_{e}c^{2}}\right)^{-p+2}\right]+\frac{1}{2-p_{\rm tail}}\left[\left(\frac{4\sigma}{\gamma_{\rm rad}}\right)^{p_{\rm tail}-2}-1\right]},$ (A4)
@@
$E_{\gamma}L_{E_{\gamma}}\approx L_{\rm max}\begin{cases}\left(\frac{E_{\gamma}}{E_{\gamma,\sigma}}\right)^{(2-p)/2}\ (E_{\gamma}<E_{\gamma,\sigma})\\ \left(\frac{E_{\gamma}}{E_{\gamma,\sigma}}\right)^{(2-p_{\rm tail})/2}\ (E_{\gamma,\sigma}<E_{\gamma}<E_{\gamma,\gamma_{\rm rad}})\\ \end{cases},$ (A5)
@@
$n_{\pm,\rm tot}$ $=\int_{E_{\rm thr}}^{E_{\gamma,\rm rad}}\frac{3L_{\rm max}^{2}\sigma_{\gamma\gamma}}{4\pi l_{\rm rec}^{3}c^{2}E_{\gamma,\sigma}^{2}}\left(\frac{E_{\gamma,\sigma}}{2m_{e}c^{2}}\right)^{p_{\rm tail}}d\log E_{\gamma,1}$ (A6)
$=\frac{3L_{\rm max}^{2}\sigma_{\gamma\gamma}}{4\pi l_{\rm rec}^{3}c^{2}E_{\gamma,\sigma}^{2}}\left(\frac{E_{\gamma,\sigma}}{2m_{e}c^{2}}\right)^{p_{\rm tail}}\log\left(\frac{E_{\gamma,\rm rad}}{E_{\rm thr}}\right),$ (A7)
@@
$n_{\pm,\rm tot}$ $=\frac{3L_{\rm max}^{2}\sigma_{\gamma\gamma}}{4\pi l_{\rm rec}^{3}c^{2}}\left[\int_{E_{\rm thr}}^{E_{\gamma,c}}\frac{1}{E_{\gamma,\sigma}^{2}}\left(\frac{E_{\gamma,\sigma}}{2m_{e}c^{2}}\right)^{p_{\rm tail}}d\log E_{\gamma,1}+\int_{E_{\gamma,c}}^{E_{\gamma,\rm rad}}E_{\gamma,\sigma}^{(p+p_{\rm tail})/2-2}(2m_{e}c^{2})^{-p}E_{\gamma,1}^{(p-p_{\rm tail})/2}d\log E_{\gamma,1}\right]$ (A8)
$=\frac{3L_{\rm max}^{2}\sigma_{\gamma\gamma}}{4\pi l_{\rm rec}^{3}c^{2}}\left[\frac{1}{E_{\gamma,\sigma}^{2}}\left(\frac{E_{\gamma,\sigma}}{2m_{e}c^{2}}\right)^{p_{\rm tail}}\log\left(\frac{E_{\gamma,\rm rad}}{E_{\rm thr}}\right)+E_{\gamma,\sigma}^{(p+p_{\rm tail})/2-2}(2m_{e}c^{2})^{-p}\frac{2}{p-p_{\rm tail}}\left(E_{\gamma,\rm rad}^{(p-p_{\rm tail})/2}-E_{\gamma,c}^{(p-p_{\rm tail})/2}\right)\right].$ (A9)
@@
$n_{\pm,\rm tot}$ $=\frac{3L_{\rm max}^{2}\sigma_{\gamma\gamma}}{4\pi l_{\rm rec}^{3}c^{2}}\left[\int_{E_{\rm thr}}^{E_{\gamma,\sigma}}\frac{1}{(2m_{e}c^{2})^{3}}\left(\frac{E_{\gamma,\sigma}}{2m_{e}c^{2}}\right)^{p-2}d\log E_{\gamma,1}+\int_{E_{\gamma,\sigma}}^{E_{\gamma,\rm rad}}E_{\gamma,\sigma}^{(p+p_{\rm tail})/2-2}(2m_{e}c^{2})^{-p}E_{\gamma,1}^{(p-p_{\rm tail})/2}d\log E_{\gamma,1}\right]$ (A10)
$=\frac{3L_{\rm max}^{2}\sigma_{\gamma\gamma}}{4\pi l_{\rm rec}^{3}c^{2}}\left[\frac{1}{(2m_{e}c^{2})^{3}}\left(\frac{E_{\gamma,\sigma}}{2m_{e}c^{2}}\right)^{p-2}\log\left(\frac{E_{\gamma,\sigma}}{E_{\rm thr}}\right)+E_{\gamma,\sigma}^{(p+p_{\rm tail})/2-2}(2m_{e}c^{2})^{-p}\frac{2}{p-p_{\rm tail}}\left(E_{\gamma,\rm rad}^{(p-p_{\rm tail})/2}-E_{\gamma,\sigma}^{(p-p_{\rm tail})/2}\right)\right].$ (A11)
Since the pair number density depends on the magnetization parameter, we calculate the magnetization parameter by solving the $\sigma_{\pm}=B_{\rm rec}^{2}/(4\pi n_{\pm,\rm tot}m_{e}c^{2})$ with Equations (A7),(A9), and (A11). We show $\sigma_{\pm}$ and $B_{\rm rec}^{2}/(4\pi n_{\pm,\rm tot}m_{e}c^{2})$ in Figure 8 as the red line and blue dashed line, respectively. We estimate the magnetization parameter as $\sigma_{\pm}\simeq 8.8\times 10^{4}$ and $1.0\times 10^{5}$ for $p=1.0,~{}p_{\rm tail}=2.5$ and $p=1.0,~{}p_{\rm tail}=3.0$. As shown in Figure 6, the estimated magnetization parameter is almost the same as estimated in the way of Chen et al. (2023) and hardly changes its value with the tail index due to self-regulation of the synchrotron spectrum to peak at $E_{\gamma}\sim m_{e}c^{2}$.

### v2

@@ Appendix E Magnetization parameter for power-law tail distribution
$\dot{N}_{e,\rm inj}=\dot{N}_{0}\begin{cases}\left(\frac{\gamma}{4\sigma}\right)^{-p}\ (\gamma<4\sigma)\\ \left(\frac{\gamma}{4\sigma}\right)^{-p_{\rm tail}}\ (4\sigma<\gamma<\gamma_{\rm rad})\\ \end{cases}.$ (E1)
@@
$L_{\rm rec}$ $\approx 2l_{\rm rec}^{2}\frac{B_{\rm rec}^{2}}{8\pi}\beta_{\rm rec}c$ (E2)
$\simeq 7.9\times 10^{41}M_{9}\dot{m}_{-4}\beta_{\rm rec,-1}\Phi_{\rm rec,1.2}^{2}\ \rm erg\ s^{-1}.$ (E3)
@@
$L_{\rm max}=\frac{L_{\rm rec}}{\frac{1}{2-p}\left[1-\left(\frac{E_{\sigma}}{m_{e}c^{2}}\right)^{-p+2}\right]+\frac{1}{2-p_{\rm tail}}\left[\left(\frac{4\sigma}{\gamma_{\rm rad}}\right)^{p_{\rm tail}-2}-1\right]},$ (E4)
@@
$E_{\gamma}L_{E_{\gamma}}\approx L_{\rm max}\begin{cases}\left(\frac{E_{\gamma}}{E_{\gamma,\sigma}}\right)^{(2-p)/2}\ (E_{\gamma}<E_{\gamma,\sigma})\\ \left(\frac{E_{\gamma}}{E_{\gamma,\sigma}}\right)^{(2-p_{\rm tail})/2}\ (E_{\gamma,\sigma}<E_{\gamma}<E_{\gamma,\gamma_{\rm rad}})\\ \end{cases},$ (E5)
@@
$n_{\pm,\rm tot}$ $=\int_{E_{\rm thr}}^{E_{\gamma,\rm rad}}\frac{3L_{\rm max}^{2}\sigma_{\gamma\gamma}}{4\pi l_{\rm rec}^{3}c^{2}E_{\gamma,\sigma}^{2}}\left(\frac{E_{\gamma,\sigma}}{2m_{e}c^{2}}\right)^{p_{\rm tail}}d\log E_{\gamma,1}$ (E6)
$=\frac{3L_{\rm max}^{2}\sigma_{\gamma\gamma}}{4\pi l_{\rm rec}^{3}c^{2}E_{\gamma,\sigma}^{2}}\left(\frac{E_{\gamma,\sigma}}{2m_{e}c^{2}}\right)^{p_{\rm tail}}\log\left(\frac{E_{\gamma,\rm rad}}{E_{\rm thr}}\right),$ (E7)
@@
$n_{\pm,\rm tot}$ $=\frac{3L_{\rm max}^{2}\sigma_{\gamma\gamma}}{4\pi l_{\rm rec}^{3}c^{2}}\left[\int_{E_{\rm thr}}^{E_{\gamma,c}}\frac{1}{E_{\gamma,\sigma}^{2}}\left(\frac{E_{\gamma,\sigma}}{2m_{e}c^{2}}\right)^{p_{\rm tail}}d\log E_{\gamma,1}+\int_{E_{\gamma,c}}^{E_{\gamma,\rm rad}}E_{\gamma,\sigma}^{(p+p_{\rm tail})/2-2}(2m_{e}c^{2})^{-p}E_{\gamma,1}^{(p-p_{\rm tail})/2}d\log E_{\gamma,1}\right]$ (E8)
$=\frac{3L_{\rm max}^{2}\sigma_{\gamma\gamma}}{4\pi l_{\rm rec}^{3}c^{2}}\left[\frac{1}{E_{\gamma,\sigma}^{2}}\left(\frac{E_{\gamma,\sigma}}{2m_{e}c^{2}}\right)^{p_{\rm tail}}\log\left(\frac{E_{\gamma,\rm rad}}{E_{\rm thr}}\right)+E_{\gamma,\sigma}^{(p+p_{\rm tail})/2-2}(2m_{e}c^{2})^{-p}\frac{2}{p-p_{\rm tail}}\left(E_{\gamma,\rm rad}^{(p-p_{\rm tail})/2}-E_{\gamma,c}^{(p-p_{\rm tail})/2}\right)\right].$ (E9)
@@
$n_{\pm,\rm tot}$ $=\frac{3L_{\rm max}^{2}\sigma_{\gamma\gamma}}{4\pi l_{\rm rec}^{3}c^{2}}\left[\int_{E_{\rm thr}}^{E_{\gamma,\sigma}}\frac{1}{(2m_{e}c^{2})^{3}}\left(\frac{E_{\gamma,\sigma}}{2m_{e}c^{2}}\right)^{p-2}d\log E_{\gamma,1}+\int_{E_{\gamma,\sigma}}^{E_{\gamma,\rm rad}}E_{\gamma,\sigma}^{(p+p_{\rm tail})/2-2}(2m_{e}c^{2})^{-p}E_{\gamma,1}^{(p-p_{\rm tail})/2}d\log E_{\gamma,1}\right]$ (E10)
$=\frac{3L_{\rm max}^{2}\sigma_{\gamma\gamma}}{4\pi l_{\rm rec}^{3}c^{2}}\left[\frac{1}{(2m_{e}c^{2})^{3}}\left(\frac{E_{\gamma,\sigma}}{2m_{e}c^{2}}\right)^{p-2}\log\left(\frac{E_{\gamma,\sigma}}{E_{\rm thr}}\right)+E_{\gamma,\sigma}^{(p+p_{\rm tail})/2-2}(2m_{e}c^{2})^{-p}\frac{2}{p-p_{\rm tail}}\left(E_{\gamma,\rm rad}^{(p-p_{\rm tail})/2}-E_{\gamma,\sigma}^{(p-p_{\rm tail})/2}\right)\right].$ (E11)
Since the pair number density depends on the magnetization parameter, we calculate the magnetization parameter by solving the $\sigma_{\pm}=B_{\rm rec}^{2}/(4\pi n_{\pm,\rm tot}m_{e}c^{2})$ with Equations (E7),(E9), and (E11). We show $\sigma_{\pm}$ and $B_{\rm rec}^{2}/(4\pi n_{\pm,\rm tot}m_{e}c^{2})$ in Figure 10 as the red line and blue dashed line, respectively. We estimate the magnetization parameter as $\sigma_{\pm}\simeq 8.8\times 10^{4}$ and $1.0\times 10^{5}$ for $p=1.0,~{}p_{\rm tail}=2.5$ and $p=1.0,~{}p_{\rm tail}=3.0$. As shown in Figure 6, the estimated magnetization parameter is almost the same as estimated in the way of Chen et al. (2023) and hardly changes its value with the tail index due to self-regulation of the synchrotron spectrum to peak at $E_{\gamma}\sim m_{e}c^{2}$.